Problem 4. 13 pts] You buy 1000 shares of Rocketman Co. for $70 each, 200 shares of Trumpology, Ind. for $50 each, and 4000 shares of Zootopia World Co. for S10 each. What are the weights in your portfolio? (Show your calculations)
Answer #1

Total value of Rocketman Co=(1000*70)=$70,000

Total value of Trumpology Inc.=(200*50)=$10,000

Total value of Zootopia World Co.=(4000*10)=$40,000

Total value=(70,000+10,000+40,000)=$120,000

Hence weight of:
Rocketman Co=(70,000/120,000)=0.5833(or 58.33% approx).

Trumpology Inc.=(10,000/120,000)=0.0833(or 8.33% approx).

Zootopia World Co.=(40,000/120,000)=0.3333(or 33.33% approx).